OpenStudy (anonymous):

Find The equestion in slope intercept form for the line contaning the two points. (1,1 ) and ( 5,4 )

OpenStudy (anonymous):

y2 - y1 slope = ------ x2 - x1 slope = 4 - 1 ----- 5-1 slope = 3/5 y = mx + b chooose any co-ordinate 4 =(3/5)(5) + b 4 = 15/3 + b 4 = 5 + b 4 - 5 = b -1 = b equation of the slope is: y = 3/5x -1

OpenStudy (anonymous):

5-1 IS 4

OpenStudy (anonymous):

How would you do this one ( 8,6 ) and (- 3,6 )